What is the electrostatic force of attraction between a 5.2 X 10 4 C charge and a 7.1 X 10 -2 C charge if they are separated by a distance of 7 meters (m)?


1
Expert's answer
2022-01-26T15:31:29-0500
"F=\dfrac 1{4\pi\epsilon_0}\dfrac {q_1q_2}{r^2}=6772\text{ N}."